A. Consider a mutual fund with $206 million in assets at the start of the year...

A.

Consider a mutual fund with $206 million in assets at the start of the year and with 20 million shares outstanding. The fund invests in a portfolio of stocks that provides dividend income at the end of the year of $2 million. The stocks included in the fund's portfolio increase in price by 6%, but no securities are sold, and there are no capital gains distributions. The fund charges 12b-1 fees of 0.50%, which are deducted from portfolio assets at year-end.


What is net asset value at the start and end of the year? (Enter your answers in dollars rounded to 3 decimal places.)

What is the rate of return for an investor in the fund?

B. You sell 490 shares of stock short that are priced at $54.15 a share. You post the 50% margin required. If the maintenance margin requirement (MMR) is 30% at what stock price do you get a margin call?
Answer #1

A). NAV0= $206,000,000/20,000,000 = $10.3

Dividends per share = $2,000,000/20,000,000 = $0.10

NAV1is based on the 6% price gain, less the 0.5% 12b-1 fee:

NAV1= $10.3 x 1.06 x (10.005) = $10.863

Rate of return = ($10.863 - $10.3 + $0.10) / $10.3 = 0.0644 = 6.44%

B). (your equity)/(value of stock owned) = 0.30

[(490)($54.15) + (0.50)(490)($54.15) - 490P]/490P = 0.30

[$26,533.50 + $13,266.75 - 490P]/490P = 0.30

$39,800.25 - 490P = 147P

$39,800.25 = 637P

P = $62.48